OpenStudy (anonymous):

Could you please show me how to solve this? 2x-1=-9

OpenStudy (anonymous):

What are you solving for?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Good, now to do that. You have to isolate it

OpenStudy (anonymous):

Meaning have it by itself

OpenStudy (anonymous):

So how would you move the 1 to the other side?

OpenStudy (anonymous):

Uhm. I do not know.

OpenStudy (anonymous):

To move it to the other side, carry the opposite operation. So add the one on both sides.

OpenStudy (anonymous):

\[2x-1=9, 2x=9+1, 2x=10, x=\frac{ 10 }{ 2}\] =5

OpenStudy (anonymous):

Do you understand that^

OpenStudy (anonymous):

No!

OpenStudy (anonymous):

@lillers4126 you get it now?

OpenStudy (anonymous):

Sensuelle, you answer is wrong

OpenStudy (anonymous):

Wouldn't it be -4? I'm probably wrong but that's what I got.

OpenStudy (anonymous):

that's correct

OpenStudy (anonymous):

Do you understand how you got it?

OpenStudy (anonymous):

Yes, I do now. I just needed help getting it started. Thank you!

OpenStudy (anonymous):

Oh okay, no problem. Remember, everyone here is going to give you answers. Don't ask for answers. Ask for help understanding.

OpenStudy (anonymous):

@OrthodoxMan my answer is not wrong if you are solving for x you get x=5, look at my steps.

OpenStudy (anonymous):

You answer is wrong. It is -9 not 9

OpenStudy (anonymous):

oaky i see it now it's -9. okay I didnt see the sign you are right in that case x=-4

OpenStudy (anonymous):

EXAMPLE NOT YOUR MATH PROBLEM y + 2 = 4 ^ isolate the variable y + 2 on the left of the = sign you want to make it just "y" how is this done y+2 + (-2) = 4 + (-2) add a -2 on both sides of the equation (whatever is done on one side of the equation must be done on the other side of the = sign) y = 2 answer